You are designing a rectangular enclosure with 3 rectangular interior sections separated by parallel walls. If you have 1500 feet of fencing, what is the maximum area that can be enclosed?
Answer by TimothyLamb(4379)

---
x = length
y = width
---
2x + 2y = 1500
a = xy
---
2x + 2y = 1500
2y = 1500 - 2x
y = (1500 - 2x)/2
y = 750 - x
---
a = xy
a = x(750 - x)
a = 750x - xx
---
answer A:
a = -xx + 750x
---
the above quadratic equation is in standard form, with a=-1, b=750, and c=0
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
-1 750 0
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic vertex is a maximum at: ( x= 375, a= 140625 )
---
the maximum area = 140,625 sq.ft
x = length = 375 ft
y = width = 375 ft
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php